# Scaling Laws for Online Advertisement Retrieval

### v1

@@ Abstract.
The scaling law is a notable property of neural network models that researchers have discovered, which greatly propels the development of large language models. Scaling laws hold great promise in guiding model design and resource allocation. Recent research increasingly shows that scaling laws are not limited to NLP tasks or Transformer architectures; they also apply to domains such as recommendation, where there are scaling laws relating to the amount of training data, model size, and model performance metrics. However, there is still a lack of literature on scaling law research in industrial online advertisement retrieval systems. This may be because 1) identifying the scaling law for resource cost and online revenue is often expensive in both time and training resources for large-scale industrial applications, and 2) varying settings for different systems prevent the scaling law from being applied across various scenarios. To address these issues, we propose a lightweight paradigm to identify the scaling law of online revenue and machine cost for a certain online advertisement retrieval scenario with a low experimental cost. Specifically, we focus on a sole factor (FLOPs) and propose an offline metric named $R/R^{*}$ that exhibits a high linear correlation with online revenue for retrieval models. We estimate the machine cost offline via a simulation algorithm. Thus, we can transform most online experiments into low-cost offline experiments. We conduct comprehensive experiments to verify the effectiveness of our proposed metric $R/R^{*}$ and to identify the scaling law in the online advertisement retrieval system of Kuaishou. With the scaling law, we demonstrate practical applications for ROI-constrained model designing and multi-scenario resource allocation in the Kuaishou advertising system. To the best of our knowledge, this is the first work to study the scaling laws for online advertisement retrieval of real-world large-scale systems, showing great potential for scaling law in advertising system optimization.
Scaling Law, Online Advertisement Retrieval, Resource Allocation
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Machine learning
1. Introduction
@@
The neural scaling laws, describing how neural network performance changes with key factors (e.g. model size, dataset size, computational cost), have been discovered in various research areas (Kaplan et al., 2020; Hoffmann et al., 2022; Shin et al., 2023; Isik et al., 2024; Zhang et al., 2024a; Fang et al., 2024). Early research (Hestness et al., 2017) shows that the neural network performance is predictable when scaling training data size in various tasks such as neural machine translation and language modeling. Kaplan et al.((2020)(Kaplan et al., 2020) further empirically verify the scaling law of Transformer architecture in language modeling, regarding the key factors (model size, data size, training cost) and training performance (PPL). Inspired by the scaling law, researchers extend the size of pre-trained language models and further empirically verify the scaling law by training GPT-3 (Brown, 2020). This wave of enthusiasm has led to the creation of GPT-3.5 and GPT-4 (Achiam et al., 2023), ushering in a new era of NLP research and applications.
@@
Based on the scaling laws, the optimal key factors of the model can be determined under given constraints, thus guiding us in model design and resource allocation. Recommendation and advertising systems are mature commercial applications that prioritize ROI, making it highly valuable to explore whether there exists scaling laws for recommendation and advertising models. Due to the lack of a thriving community and open data ecosystem similar to NLP, research on model scaling is relatively scarce in the recommendation and advertisement areas. Early studies primarily gave some qualitative conclusions about model scaling (Shin et al., 2023; Zhang et al., 2024a). Fang et al.((2024)(Fang et al., 2024) first attempt to give a quantitative scaling law of model performance and the amount of training data and model parameters based on public information retrieval benchmarks, and give the practice to solve the optimal amount of data and model parameters under a given total training resource.
@@
However, there is still a lack of literature on scaling law research in real-world online recommendation and advertising systems. We attribute this to two main challenges: 1) For commercial systems, the scaling law should describe the relationship between business revenue and machine costs, rather than the relationship between computing volume, data volume, and offline metrics as seen in traditional research. Achieving this requires training numerous models and conducting extensive online experiments, which are prohibitively expensive in terms of time, manpower, and machine costs; and 2) varying settings across different systems prevent the scaling law from being universally applicable. As a result, to use scaling laws to guide system optimization in a specific scenario, one must first incur substantial costs to obtain the necessary parameters. This makes it impractical for commercial systems to identify and apply scaling laws in many real-world settings.
@@
We aim to identify scaling laws in online advertising systems with low experimental costs and explore their applications in guiding system optimization. An advertising system typically consists of several subsystems, including the advertisement retrieval subsystem and the bidding and ranking subsystem. Each subsystem has its own technology stack, making it challenging to establish a unified paradigm for lightweight identification of scaling laws. In this work, we focus on the online advertisement retrieval subsystem, which selects the top-k ads for the next stage but does not determine the billing of ads. This subsystem is described in detail in Section 3.
To address the issue of high experimental costs, we propose a lightweight paradigm for identifying scaling laws in online advertisement retrieval. First, we introduce a novel offline metric, $R/R^{*}$, which incorporates the revenue information of each ad in the training data and exhibits a high linear correlation with online revenue for retrieval models. Note that we can obtain the specific relationship between $R/R^{*}$ and online revenue for our scenario through the records of daily iterations, theoretically eliminating the need for additional online experiments. Next, we conduct extensive offline experiments to verify the scaling behavior of MLP models in the retrieval scenario, where FLOPs serve as the independent variable and $R/R^{*}$ serves as the dependent variable. Our results show that these scaling behaviors follow a broken neural scaling law (Caballero et al., 2023). Finally, we propose a simple simulation algorithm, developed by our engineering team, to estimate the machine cost. By using the model size parameter as the decision variable and FLOPs and $R/R^{*}$ as intermediate variables, we can associate the model’s estimated machine cost with the estimated online revenue.
We conduct extensive online experiments to verify the effectiveness of our proposed $R/R^{*}$ metric. The results show that $R/R^{*}$ is a more advanced offline surrogate metric for online revenue compared to traditional metrics. When there is no revenue-related information recorded at the ad granularity in the training data, $NDCG$ can serve as a suitable substitute for $R/R^{*}$. Additionally, we demonstrate the applications of the scaling law in ROI-constrained model design and multi-scenario resource allocation. These applications result in a 0.85% and 2.8% improvement in online revenue, respectively, highlighting the significant value of the scaling law in guiding system optimization.
In general, our main contributions are three-fold: 1) We introduce an offline metric, $R/R^{*}$, which exhibits a strong linear correlation with online revenue, bringing significant convenience to the validation of scaling laws and the iterative development of retrieval models. 2) We provide a lightweight paradigm to identify the scaling laws between machine cost and online revenue and verify the existence of scaling laws in industrial advertisement retrieval scenarios. 3) We demonstrate practical applications of the scaling laws in the Kuaishou advertising system. With the scaling law of online revenue and machine cost, we can efficiently perform offline ROI estimations for hundreds of models with different sizes using just a few machines within a few days. The models obtained through this process have led to significant revenue improvement, highlighting the potential of scaling laws in optimizing advertising systems.
2. Related Work
2.1. Neural Scaling Laws
The neural scaling laws, widely recognized in Natural Language Processing and Computer Vision areas, establish predictable relationships between model performance and key factors such as model size, dataset size, computational cost, and post-training error rates. Hestness et al.((2017)(Hestness et al., 2017) introduced a unified scaling law applicable to machine translation, image classification, and speech recognition, noting that different tasks exhibit distinct scaling coefficients. Kaplan et al.((2020)(Kaplan et al., 2020) further elaborated on these laws, defining them through four parameters: model size (number of parameters), dataset size (number of tokens), computational cost (FLOPs), and training loss (perplexity per token). These relationships were empirically validated, including during the training of GPT-3 (Brown, 2020). Subsequently, Hoffmann et al.((2022)(Hoffmann et al., 2022) presented the Chinchilla Scaling Law, which differs somewhat from (Kaplan et al., 2020) because of their different training setups. In addition to estimating the training loss of the model, Isik et al.((2024)(Isik et al., 2024) further verified that there also exist scaling laws between the downstream task metrics and the model parameters. Caballero et al.((2023)(Caballero et al., 2023) found that many scaling behaviors of artificial neural networks follow a smoothly broken power law (BNSL), verified on various NLP and CV tasks, covering a wide range of downstream tasks and model structures.
In the recommendation area, Shin et al.((2023)(Shin et al., 2023) and Zhang et al.((2024a)(Zhang et al., 2024a) conducted studies on whether there exist scaling laws for recommendation models and primarily provided qualitative conclusions. Fang et al.((2024)(Fang et al., 2024) first proposed a quantitative scaling law in the recommendation area, which describes the relationship between the amount of training data, the size of model parameters, and an offline metric for query-document retrieval. Based on this scaling law, the optimal amount of data and model parameter allocation can be solved under a given total training resource. However, obtaining a multivariate scaling law requires a large number of experiments, and the offline metrics might not be a good indicator for online metrics, these make it somewhat difficult to apply in real-world industrial applications. In this paper, we focus on how to obtain the scaling law function between the model’s scalable hyper-parameters (such as the FLOPs) and the online revenue (the primary goal of the advertising system) with only a small amount of experimental cost.
2.2. Models and Evaluation for Online Advertisement Retrieval
Online advertising systems often adopt a cascade ranking framework (Qin et al., 2022; Wang et al., 2011; Gallagher et al., 2019; Chen et al., 2017). The cascade ranking usually includes two types of stages, namely retrieval and ranking. The retrieval stages take a set of terms as input and supply the top-k predicted terms to the next stage, which mainly focuses on order accuracy. The ranking stages in advertising systems should focus on not only the order accuracy but also the calibration accuracy (Huang et al., 2022; Sheng et al., 2023). These lead to some technique differences in retrieval and ranking stages, such as optimization objectives, surrogate losses, design of evaluation metrics, etc. In this work, we focus on the retrieval stages of online advertising systems.
The retrieval stages select the top-k set for the next stage, typically including multiple ranking pathways. In the retrieval stages, which normally refers to the ”Matching” or ”Pre-rank” stage in industrial systems, models often adopt a twin-tower (Covington et al., 2016) or MLP (Hornik et al., 1989) architecture. Some systems may also adopt more complex architectures like transformers (Vaswani et al., 2017). In terms of objective design, industrial systems usually aim to estimate CTR (He et al., 2014) and exposure probability of ads, and use learning-to-rank methods (Wang et al., 2018, 2024) to learn the results of ranking models. For evaluation of retrieval tasks, researchers commonly use OPA, NDCG, and Recall (Wang et al., 2024; Zangerle and Bauer, 2022; Bauer et al., 2024). However, existing metrics often have complex relationships with online revenue, thus we propose $R/R^{*}$, which exhibits a strong linear relationship with online revenue, to reduce the experimental cost of identifying the scaling law for industry scenarios.
3. Formulation of Online Advertisement Retrieval
3.1. Overview of Online Advertising System
@@
Figure 1 illustrates a typical cascade ranking system for online advertisements deployed at Kuaishou Technology. This system is comprised of four main stages: Matching, Pre-ranking, Ranking, and Mix-ranking. The ”Matching” and ”Pre-ranking” stages serve as retrieval mechanisms, focusing on selecting candidate ads without directly impacting their billing. Conversely, the ”Ranking” stage plays a dual role by determining both the selection and billing of ads. It achieves this by predicting the eCPM (Effective Cost Per Mille) of each ad, serving as the standard for billing after its exposure. The ”Mix-ranking” stage integrates the outputs from the advertising and recommendation systems to decide the final set of items presented to the user. For advertising systems, the differences in the goal of the retrieval and ranking stages lead to some technical differences, such as training objectives, system design, evaluation metrics, etc. In this work, we focus on the scaling laws of retrieval stages.
@@
Figure 2 shows the algorithm architecture of the Matching and Pre-ranking stages in the Kuaishou advertising system. They all adopt a multi-pathway ensemble architecture. The Matching stage can be divided into model-based pathways and rule-based pathways. The rule-based pathway can quickly filter out irrelevant ads using predefined criteria, ensuring compliance with regulations and improving ad relevance, while providing flexibility and transparency in the ad selection process. The model-based pathway uses machine learning algorithms to predict and select ads, achieving higher performance ceilings in ad selection. The Pre-ranking stage operates on the set of ads already selected by the Matching stage, and thus only employs the model-based pathways.
@@
Model-based pathways in the Matching and Pre-ranking stages can typically be categorized based on different modeling objectives, such as revenue-oriented, user-interest-oriented, and lifetime value-oriented pathways. In advertising systems, revenue-oriented pathways usually carry the most weight. Since the earlier stages only need to perform set selection, Kuaishou’s advertising system employs Learning-to-Rank (LTR) methods to learn the optimal ranking order to maximize the revenue objective. The LTR pathway is the most prevalent and important in the system, as it is typically the primary focus for algorithm researchers to iterate and improve. Therefore, we use the Pre-ranking LTR pathway as a case study to verify scaling laws in online advertisement retrieval, which similarly applies to Matching models because they share the same multi-pathway architectures.
@@
Figure 1. The cascade ranking architecture at Kuaishou Technology, includes four stages: Matching, Pre-ranking, Ranking, and Mix-ranking. Typically, we regard the ”Matching” and ”Pre-ranking” stages as retrieval stages, which do not decide the billing of ads.
@@
Figure 2. The multi-pathway architecture of Matching and Pre-ranking stages in Kuaishou advertising system. Each pathway has its own weight, which determines the proportion of ads sent to the next stage after the Fusion module.
3.2. Formulation of Retrieval Tasks
@@
Here we give the formulation of retrieval tasks represented by learning-to-rank in the Pre-ranking stage. The organization of training samples and the surrogate loss follows the ARF (Wang et al., 2024). We randomly draw samples from the Pre-ranking, Ranking, and Mix-ranking stages. The training set can be formulated as:
(1) $D_{train}=\{(f_{u_{i}},\{f_{a_{i}^{j}},v_{i}^{j},ecpm_{i}^{j}|1\leq j\leq n\})_{i}\}_{i=1}^{N}$
@@
where $u_{i}$ means the user of the $i$-th impression in the training set, $a^{j}$ means the $j$-th material for ranking in the system. $N$ is the number of impressions of $D_{train}$. The $i$ in $a_{i}^{j}$ means that the sample $a^{j}$ corresponds to impression $i$. The size of the materials for each impression is $n$. $f_{(\cdot)}$ means the feature of ()($\cdot$)( ⋅ ). $v_{i}^{j}$ is the ground truth rank index (the higher value is considered better) of the pair $(u_{i},a^{j})$. The rank index is the relative position of the sampled ads within the system queue, ordered by their original positions. $ecpm_{i}^{j}$ is the eCPM (Effective Cost Per Mille) predicted by the Ranking stage, which can be regarded as the expected revenue for the exposure of $ad_{i}^{j}$. If $ad_{i}^{j}$ does not win in the Pre-ranking and thus does not enter the Ranking stage, the $ecpm_{i}^{j}$ equals 0. The eCPM information in the dataset is only used to construct our offline evaluation metric.
Our learning target is to let the model pick up all the ground-truth ads to the next stage. The surrogate loss of the ARF (Wang et al., 2024) can be formulated as follows:
(2) $L_{total}=L_{Relax}+\frac{1}{2\alpha^{2}}L_{Global}+\log(|\alpha|)$
In Eq 2, $\alpha$ is a trainable scalar, the $L_{Relax}$ and $L_{Global}$ can be formulated as Eq 3 and Eq 4 in the following:
(3) $\begin{split}L_{Relax}&=-\sum_{j}^{n}\{[\sum^{col}\hat{\mathcal{P}}^{\downarrow}_{\textbf{v}_{i}}[:k]]\circ[\log(\frac{1}{m}\sum^{col}\hat{\mathcal{P}}^{\downarrow}_{\mathcal{M}^{(i,\cdot)}}[:m])]\}_{j}\end{split}$
(4) $\begin{split}L_{Gobal}&=-\sum_{j}^{n}\sum_{h}^{n}[\hat{\mathcal{P}}^{\downarrow}_{\textbf{v}_{i}}[j,:]\circ\log(\hat{\mathcal{P}}^{\downarrow}_{\mathcal{M}^{(i,\cdot)}}[j,:])]_{h}\end{split}$
where $\hat{\mathcal{P}}^{\downarrow}_{(\cdot)}$ denote the soft permutation matrices obtained by a differentiable sorting operator (e.g. NeuralSort (Grover et al., 2019)) for sorting vector ()($\cdot$)( ⋅ ) in descending order, $\mathcal{M}^{(i,\cdot)}$ denotes the model predicted score vector of the i-th impression, $\textbf{v}_{i}$ denotes the vector of the ground-truth value for the ads in the i-th impression. $\circ$ refers to the element-wise matrix multiplication operation. $\sum^{col}$ is the column sum for a 2-D matrix. {}$\{\cdot\}_{t}$ and $[\cdot]_{t}$ in Eq 3 and Eq 4 denote the t-th element of the tensor $\cdot$. $m$ and $k$ are hyper-parameters of the ARF.
3.3. Features and Foundation Model
Regarding the features, we adopt both sparse and dense features to describe the information of users and ads in the online advertising system. Sparse features are those whose embeddings are obtained from embedding lookup tables, while dense features are the raw values themselves. The sparse features of the user primarily include the action list of ads and user profile information (e.g., age, gender, and region). The action list mainly consists of action types, frequencies, target ads, and timestamps. The sparse features of the ads primarily include the IDs of the ad and its advertiser. The user’s dense features mainly consist of embeddings produced by other pre-trained models. The dense features of the ads primarily include statistical features information and multimodal features generated by multimodal models.
Regarding the foundation model, we employ the multi-layer perceptrons (MLPs) architecture. The MLP consists of 5 layers, where each hidden layer is composed of batch normalization, linear mapping, and a PReLU (He et al., 2015) activation function in sequence. The output layer is a pure linear layer. Parameters are initialized using the He initialization (He et al., 2015). We apply a log1p transformation as in (Qin et al., 2021) for all statistical dense features. All sparse and dense features are concatenated together and fed as input to the model.
3.4. Background of Online Training and Serving
Our model is trained online using a distributed training framework with synchronous training. We use Adagrad (Duchi et al., 2011) as the optimizer with a learning rate of 0.01 and an epsilon value of 1e-8. The training task consumes data in a streaming manner, approximately processing samples in the order they are generated. The data stream ensures that ads within the same impression are aggregated together and perform a shuffle at the impression level within a small time interval. We use single precision floating point format (FP32) for training and saving the model.
We use a TensorRT-based service architecture to deploy the model for online serving. To accelerate the computation during inference, we adopt half-precision floating point format (FP16) for each layer. Additionally, we decompose the first-layer computation of the MLP into separate user-side and item-side calculations, which are then summed using TensorFlow’s broadcasting mechanism. In this way, the user-side computation needs to be performed only once per impression, while the first-layer item-side computation is cached and periodically refreshed by a dedicated service. We named this strategy ”first-layer optimization”. Since most of the computational cost of an MLP is concentrated in the first layer, the ”first-layer optimization” significantly reduces this cost compared to the original calculation strategy. However, it also introduces a more complex relationship between the model’s FLOPs and the associated machine costs. Due to space limitations, more details about online training and serving are described in appendix A.
4. Scaling Laws of Budget and Revenue

### v2

@@ Abstract
The scaling law is a notable property of neural network models and has significantly propelled the development of large language models. Scaling laws hold great promise in guiding model design and resource allocation. Recent research increasingly shows that scaling laws are not limited to NLP tasks or Transformer architectures; they also apply to domains such as recommendation. However, there is still a lack of literature on scaling law research in online advertisement retrieval systems. This may be because 1) identifying the scaling law for resource cost and online revenue is often expensive in both time and training resources for industrial applications, and 2) varying settings for different systems prevent the scaling law from being applied across various scenarios. To address these issues, we propose a lightweight paradigm to identify online scaling laws of retrieval models, incorporating a novel offline metric $R/R^{*}$ and an offline simulation algorithm. We prove that under mild assumptions, the correlation between $R/R^{*}$ and online revenue asymptotically approaches 1 and empirically validates its effectiveness. The simulation algorithm can estimate the machine cost offline. Based on the lightweight paradigm, we can identify online scaling laws for retrieval models almost exclusively through offline experiments, and quickly estimate machine costs and revenues for given model configurations. We further validate the existence of scaling laws across mainstream model architectures—including Transformer, MLP, and DSSM—in our real-world advertising system. With the identified scaling laws, we demonstrate practical applications for ROI-constrained model designing and multi-scenario resource allocation in the online advertising system. To the best of our knowledge, this is the first work to study the identification and application of online scaling laws for online advertisement retrieval, showing great potential for scaling laws in advertising system optimization.
1 Introduction
@@
The neural scaling laws, describing how neural network performance changes with key factors (e.g. model size, dataset size, computational cost), have been discovered in various research areas [31, 25, 43, 29, 55, 18]. Early research [24] shows that the neural network performance is predictable when scaling training data size in various tasks such as neural machine translation and language modeling. Kaplan et al.(2020) further empirically verify the scaling law of the Transformer architecture in language modeling, regarding the key factors (model size, data size, training cost) and training performance (PPL). Inspired by the scaling law, researchers extend the size of pre-trained language models and further empirically verify the scaling law by training GPT-3 [6]. This wave of enthusiasm has led to the creation of GPT-3.5 and GPT-4 [3], ushering in a new era of NLP research and applications.
@@
Based on the scaling laws, the optimal key factors of the model can be determined under given constraints, thus guiding us in model design and resource allocation. Recommendation and advertising systems are mature commercial applications that prioritize ROI, making it highly valuable to explore whether there exist scaling laws for recommendation and advertising models. Due to the lack of a thriving community and open data ecosystem similar to NLP, research on model scaling is relatively scarce in the recommendation and advertisement areas. Early studies primarily gave some qualitative or offline quantification conclusions about model scaling [43, 55]. Fang et al.(2024) attempts to give a quantitative scaling law of model performance and the amount of training data and model parameters based on public information retrieval benchmarks, and first gives an offline application practice in the recommendation area, which is to solve the optimal amount of data and model parameters under a given total training resource.
@@
Crucially, no work has addressed the online identification of scaling laws between business revenue and machine costs in real-world advertising systems. We attribute this to two primary challenges: 1) The scaling law must describe the online revenue-cost relationship (not offline metrics), requiring costly online experiments; 2) System-specific configurations prevent universal applicability, making scaling law adoption prohibitively expensive. As a result, to use scaling laws to guide system optimization in a specific scenario, one must first incur substantial costs to obtain the necessary parameters. This makes it impractical for commercial systems to identify and apply scaling laws in many real-world settings.
@@
We aim to identify scaling laws in online advertising systems with low experimental costs and explore their applications in guiding system optimization. An advertising system typically consists of several subsystems (e.g., advertisement retrieval, bidding, and ranking), each with distinct technical architectures. This heterogeneity complicates establishing a unified paradigm for lightweight scaling law identification. In this work, we focus on the online advertisement retrieval subsystem—a critical component that selects the top-k ads for downstream stages without directly determining billing. This subsystem is described in detail in Section 2.
To address these challenges, we propose a lightweight paradigm for identifying online scaling laws in industrial advertisement retrieval. First, we introduce a novel offline metric, $R/R^{*}$, which integrates the predicted revenue of each ad from the training data. We prove that, under mild assumptions, the correlation between $R/R^{*}$ and online revenue asymptotically approaches 1. This theoretical guarantee, combined with historical A/B test data from daily model iterations, enables us to calibrate the $R/R^{*}$–revenue relationship without requiring additional online experiments. Then, we design a simulation algorithm to estimate machine cost from model configurations. By treating model size as the decision variable and using FLOPs and $R/R^{*}$ as intermediate proxies, our paradigm enables end-to-end prediction of both machine cost and online revenue, facilitating efficient offline model selection and system optimization.
We validate the effectiveness of $R/R^{*}$ through extensive online A/B tests. Results demonstrate that $R/R^{*}$ serves as a more accurate offline surrogate for online revenue than traditional metrics. Then, we conduct extensive offline experiments across mainstream retrieval architectures—Transformer [45], MLP [1], and DSSM [27]—in both the pre-ranking and matching stages. Using FLOPs as the independent variable and $R/R^{*}$ as the dependent variable, we observe consistent broken neural scaling laws [9] across all architectures and stages. Building on this, we apply the identified scaling laws to two critical use cases: ROI-constrained model design and multi-scenario resource allocation. These deployments yield a substantial combined 5.10% improvement in online revenue, demonstrating the practical utility of scaling laws in real-world system optimization.
Our contributions are threefold: 1) We propose $R/R^{*}$, a novel offline metric that is both theoretically justified (asymptotically correlated with online revenue under mild assumptions) and empirically validated as a superior surrogate for online revenue. 2) We present a lightweight paradigm for identifying scaling laws between machine cost and online revenue in industrial retrieval systems, enabling offline discovery of broken neural scaling laws across diverse architectures (e.g., Transformer, MLP, DSSM). 3) We demonstrate real-world applications in ROI-constrained model selection and multi-scenario resource allocation, achieving a substantial 5.10% gain in online revenue. Our framework enables rapid offline ROI estimation for hundreds of model configurations within days—accelerating model iteration and system optimization at scale.
2 Formulation of Online Advertisement Retrieval
In this section, we clarify the research subject and experimental background of this paper. We give a formulation of online advertisement retrieval, including the definition of the retrieval stage, mainstream industrial practice of architecture and learning tasks for retrieval (i.e., multi-pathway architecture and learning-to-rank tasks).
2.1 Overview of Online Advertising System
@@
Figure 1 illustrates a typical cascade ranking system for online advertising systems. The system is comprised of four main stages: Matching, Pre-ranking, Ranking, and Mix-ranking. The “Matching” and “Pre-ranking” stages serve as retrieval mechanisms, focusing on selecting candidate ads without directly impacting their billing. Conversely, the “Ranking” stage plays a dual role by determining both the selection and billing of ads. It achieves this by predicting the eCPM (Effective Cost Per Mille) of each ad, serving as the standard for billing after its exposure. The “Mix-ranking” stage integrates the outputs from the advertising and recommendation systems to decide the final set of items presented to the user. For advertising systems, the differences in the goal of the retrieval and ranking stages lead to some technical differences, such as training objectives, system design, evaluation metrics, etc. In this work, we focus on the scaling laws of retrieval stages.
@@
Figure 2 shows the typical algorithm architecture of the Matching and Pre-ranking stages in advertising systems. They all adopt a multi-pathway ensemble architecture. The Matching stage can be divided into model-based pathways and rule-based pathways. The rule-based pathway can quickly filter out irrelevant ads using predefined criteria, ensuring compliance with regulations and improving ad relevance, while providing flexibility and transparency in the ad selection process. The model-based pathway uses machine learning algorithms to predict and select ads, achieving higher performance ceilings in ad selection. The Pre-ranking stage operates on the set of ads already selected by the Matching stage, and thus only employs the model-based pathways.
@@
Model-based pathways in the Matching and Pre-ranking stages can typically be categorized based on different modeling objectives, such as revenue-oriented, user-interest-oriented, and lifetime value-oriented pathways. In advertising systems, revenue-oriented pathways usually carry the most weight. Since the earlier stages only need to perform set selection, advertising systems often employ Learning-to-rank methods to learn the optimal ranking order to maximize the revenue objective. The Learning-to-rank pathway is the most prevalent and important in the system, as it is typically the primary focus for algorithm researchers to iterate and improve [48, 30, 57, 35, 49, 50]. Therefore, we use the Learning-to-rank pathway as a case study to verify scaling laws in online advertisement retrieval.
@@
Figure 1: The cascade ranking architecture of our advertising system, includes four stages: Matching, Pre-ranking, Ranking, and Mix-ranking. Typically, we regard the “Matching” and “Pre-ranking” stages as retrieval stages, which do not decide the billing of ads.
@@
Figure 2: A typical multi-pathway architecture of Matching and Pre-ranking stages. Each pathway has its own weight, determining the proportion of ads sent to the next stage after the Fusion module.
2.2 Formulation of Retrieval Tasks
@@
Here, we give the formulation of retrieval tasks represented by learning-to-rank. Following Zheng et al., Wang et al.(2024, 2025), we randomly draw samples from the Matching, Pre-ranking, Ranking, and Mix-ranking stages. We use all the samples for training the model of the Matching stage, and we use the samples from all stages except the Matching stage to train the Pre-ranking models. The training set can be formulated as:
$D_{train}=\{(f_{u_{i}},\{f_{a_{i}^{j}},v_{i}^{j},ecpm_{i}^{j}|1\leq j\leq n\})_{i}\}_{i=1}^{N}$ (1)
@@
where $u_{i}$ means the user of the $i$-th impression in the training set, $a^{j}$ means the $j$-th material for ranking in the system. $N$ is the number of impressions of $D_{train}$. The $i$ in $a_{i}^{j}$ means that the sample $a^{j}$ corresponds to impression $i$. The size of the materials for each impression is $n$. $f_{(\cdot)}$ means the feature of $(\cdot)$. $v_{i}^{j}$ is the ground truth rank index (the higher the better) of the pair $(u_{i},a^{j})$. The rank index is the relative position of the sampled ads within the system queue, ordered by their original positions. $ecpm_{i}^{j}$ is the eCPM (Effective Cost Per Mille) predicted by the Ranking stage, which can be regarded as the expected revenue for the exposure of $ad_{i}^{j}$. If $ad_{i}^{j}$ does not win in the Matching or Pre-ranking and thus does not enter the Ranking stage, the $ecpm_{i}^{j}$ equals 0. The eCPM information in the dataset is only used to construct our offline evaluation metric.
The learning objective in Learning-to-Rank is to learn the order of all or part of the training set $D_{train}$. This can typically be achieved using different types of methods, such as point-wise [14, 41, 12, 32], pair-wise [10, 58, 7], and list-wise [8, 48, 49, 57] approaches. We employ ARF [49] to train the models in our scenario.
3 Scaling Laws of Budget and Revenue